# Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization

### v2

@@ Abstract
Combinatorial optimization problems involving multiple agents are notoriously challenging due to their NP-hard nature and the necessity for effective agent coordination. Despite advancements in learning-based methods, existing approaches often face critical limitations, including suboptimal agent coordination, poor generalizability, and high computational latency. To address these issues, we propose Parallel AutoRegressive Combinatorial Optimization (PARCO), a reinforcement learning framework designed to construct high-quality solutions for multi-agent combinatorial tasks efficiently. To this end, PARCO integrates three key components: (1) transformer-based communication layers to enable effective agent collaboration during parallel solution construction, (2) a multiple pointer mechanism for low-latency, parallel agent decision-making, and (3) priority-based conflict handlers to resolve decision conflicts via learned priorities. We evaluate PARCO in multi-agent vehicle routing and scheduling problems where our approach outperforms state-of-the-art learning methods and demonstrates strong generalization ability and remarkable computational efficiency.
Machine Learning, ICML
@@ 1 Introduction
Combinatorial optimization (CO) problems involve determining an optimal sequence of actions in discrete spaces with several crucial domains, including logistics and supply chain management (Weinand et al., 2022). Many practical CO problems require a solution to be constructed by coordinating multiple distinct entities (i.e., agents), each with unique characteristics. We call such problems multi-agent CO. This class of problems naturally arises in real-world applications such as coordinated vehicle routing for disaster management (Liberatore et al., 2014), manufacturing (Jungwattanakit et al., 2008) and last-mile delivery optimization (Archetti & Bertazzi, 2021), where heterogeneous agents must operate under complex constraints.
@@
CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time due to their NP-hardness (Jünger et al., 1995; Peres & Castelli, 2021). While traditional methods such as exact and heuristic methods have been developed to solve a variety of problems (Laporte & Osman, 1995; Hejazi & Saghafian, 2005), these approaches often concentrate on single-agent scenarios and require long execution times. Moreover, multi-agent CO poses numerous additional challenges such as additional constraints and different optimization objectives including minimizing lateness or the makespan among agents (Soman & Patil, 2020; Mahmoudinazlou & Kwon, 2024).
@@
Recently, Neural Combinatorial Optimization (NCO) has emerged as a powerful alternative by learning efficient neural solvers (Bengio et al., 2021). In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with CO problems instead of relying on costly labeled datasets in the shape of optimal solutions, and could even outperform traditional approaches by automatically discovering better (neural) heuristics (Bello et al., 2016; Kwon et al., 2020; Kim et al., 2023a). Among NCO methods, Autoregressive (AR) models—constructing solutions step-by-step—have garnered attention for their ability to generate solutions for a variety of problems with hard constraints (Kool et al., 2019a; Kwon et al., 2021; Bi et al., 2024). This capability is crucial for addressing complex problems with multiple constraints, such as heterogeneous capacity (Baldacci et al., 2008) or precedences in pickup and delivery (Parragh et al., 2008) and machine scheduling problems (Zhang et al., 2020a).
@@
However, despite their promising results in single-agent settings, AR-CO methods pose two challenges that hinder their practical adoption in solving complex multi-agent CO problems. First, existing AR methods for multi-agent CO suffer from poor coordination, resulting in unsatisfactory solution quality and poor generalization across varying problem sizes and agent configurations (Zong et al., 2022; Son et al., 2024; Zheng et al., 2024a; Liu et al., 2024b). Moreover, AR sequence generation is associated with high latency due to each single action (or “token”) depending on each single previous one, akin to large language models (Leviathan et al., 2023; Gao et al., 2024). This issue becomes increasingly pronounced when dealing with large problem instances involving numerous agents.
Figure 1: Example multi-agent vehicle routing problem. [Top] Autoregressive policies construct solutions one agent at a time, with inefficient coordination and requiring several steps. [Bottom] Our parallel autoregressive policies ensure not only improved decision-making thanks to inter-agent communication that enables effective coordination but also faster construction in fewer steps.
@@
This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel learning framework to address multi-agent combinatorial problems effectively via parallel solution construction as illustrated in Figure 1. We design specialized transformer-based Communication Layers to enhance coordination among agents during decision-making, enabling collaborative behaviors. Our model leverages a Multiple Pointer Mechanism to efficiently generate actions for various agents simultaneously at each solution construction step. Priority-based Conflict Handlers ensure feasibility and resolve potential conflicts among agents’ decisions based on learned priorities.
@@
  • We evaluate PARCO on multi-agent vehicle routing and scheduling, where we outperform state-of-the-art learning methods across diverse problems [^1] Code at: https://github.com/ai4co/parco.
@@ 2 Related Work
Recent advancements in Neural Combinatorial Optimization (NCO) have shown promising end-to-end solutions for combinatorial optimization problems (Bengio et al., 2021; Yang & Whinston, 2023). NCO has led to the development of a variety of methods for diverse problems, including the incorporation of problem-specific biases (Jin et al., 2023; Luo et al., 2023; Kim et al., 2023c, b; Son et al., 2023; Gao et al., 2023), bi-level solution pipelines (Li et al., 2021b; Ye et al., 2023b; Zheng et al., 2024b), learning-guided search (Ye et al., 2023a; Kim et al., 2024; Sun & Yang, 2023; Li et al., 2024b, 2025), improvement methods (Hottung et al., 2021; Ma et al., 2021, 2024), effective training algorithms (Kim et al., 2022; Grinsztajn et al., 2023; Drakulic et al., 2023; Corsini et al., 2024; Pirnay & Grimm, 2024; Luo et al., 2024), downstream applications (Chen et al., 2023; Zhou et al., 2023; Ma et al., 2023; Luttmann & Xie, 2024; Yan & Wu, 2024), and the recent development of end-to-end foundation models (Liu et al., 2024a; Zhou et al., 2024; Drakulic et al., 2024; Berto et al., 2024b; Jiang et al., 2024; Li et al., 2024a). Among such methods, RL-based end-to-end autoregressive (AR) models present several advantages, including eliminating the need for labeled solutions, reducing reliance on handcrafted heuristics, and achieving high efficiency in generating high-quality solutions (Kool et al., 2019a; Kwon et al., 2020, 2021; Berto et al., 2024a; Bonnet et al., 2024).
@@
While the seminal works in AR-CO methods of Vinyals et al. (2015); Kool et al. (2019a); Kwon et al. (2020, 2021) propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes and constraints. Building on Kool et al. (2019a), Son et al. (2024) and Zheng et al. (2024a) introduce attention-based policies for multi-agent min-max routing. These models adopt a sequential AR construction strategy, solving for one agent at a time and switching agents only after completing a single-agent solution. While this approach outperforms decentralized methods (Cao et al., 2022; Park et al., 2023) it remains inherently sequential. In contrast, other multi-agent AR methods determine a location for every agent in each solution construction step, but the different agents select their actions sequentially in either random (Zhang et al., 2020b) or learned order (Falkner & Schmidt-Thieme, 2020; Li et al., 2022; Liu et al., 2024b), making these models still suffer from high generation latency as well as missing inter-agent communication and coordination. Zong et al. (2022) propose a multi-agent pickup and delivery model with parallel decoding, using distinct decoders for each agent. However, this approach exhibits limited generalizability due to inflexible fixed decoders for specific agents, lacking a powerful communication mechanism, and conflict resolution handled naively by assigning random precedence to agents, restricting the robustness of the model. PARCO aims at addressing the shortcomings of previous works by leveraging parallel solution construction for any number of agents efficiently with a Multiple Pointer Mechanism, enhancing coordination via Communication Layers, and solving conflicts in a principled manner via Priority-based Conflict handlers. Finally, unlike previous works, PARCO tackles a general framework tackling multi-agent CO without restricting to a single class of problems.
@@ 3.1 Markov Decision Processes
Multi-agent CO problems can be formulated as Markov Decision Processes (MDPs) and solved autoregressively using RL (Mazyavkina et al., 2021). In this framework, a solution $\bm{a}$ to a CO problem instance $\bm{x}$ is represented as a sequence of actions. Actions $a_{t}$ are selected sequentially from the action space $\mathcal{A}$ of size $N$ based on the current state $s_{t}\in\mathcal{S}$, which encodes the problem’s configuration at step $t$. In multi-agent problems, at each step one agent $m\in\mathcal{M}=\{1,\dots,M\}$ selects an action $a^{m}_{t}$ according to a policy $\pi_{\theta}$, usually represented by a $\theta$-parametrized neural network, mapping states to actions. Agents are selected either by some predefined precedence rule as in the sequential planning of Son et al. (2024), where agent solutions are constructed one after another, or by the policy itself, in which case $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}\times\mathcal{M}$. Given the agent and its corresponding action, the problem then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $\tau:\mathcal{S}\times\mathcal{A}\times\mathcal{M}\rightarrow\mathcal{S}$. This process reaches the terminal state once it has generated a feasible solution $\bm{a}=(a_{1},...,a_{T})$ for the problem instance $\bm{x}$ in $T$ construction steps. The (sparse) reward $R(\bm{a},\bm{x})$ is usually obtained only in the terminal state and takes the form of the negative of the cost function of the respective CO problem.
@@ 3.2 AR Models for CO
Given the sequential nature of MDPs, AR methods pose a natural choice for the policy $\pi_{\theta}$. AR methods construct a viable solution by sequentially generating actions based on the current state and previously selected actions. Without loss of generality, the process can be represented in an encoder-decoder framework as:
@@
where $\bm{a}_{<t}=(a_{1},\ldots,a_{t-1})$ is the sequence of actions taken prior to $t$ and $\bm{h}=f_{\theta}(\bm{x})$ is an encoding of the problem instance $\bm{x}$ obtained via the encoder network $f$. The decoder $g_{\theta}$ then autoregressively generates the sequence of actions, conditioned on $\bm{h}$ and the previously generated actions. The parameters $\theta$ encompass both the encoder and decoder components, which together define the policy as the mechanism for producing the joint distribution $p_{\theta}(\bm{a}|\bm{x})$.
3.3 RL Training
RL leverages the MDP formulation to learn the parameters $\theta$ of the policy $\pi_{\theta}$ without relying on labeled optimal solutions. We can thus cast the optimization problem as follows:
[…]
where $P(\bm{x})$ is the distribution over problem instances and $R(\bm{a},\bm{x})$ is the reward of a solution. We can solve the optimization problem in Equation 2 using various RL methods, including policy gradient (REINFORCE) variations (Kool et al., 2019b; Kwon et al., 2020; Kim et al., 2022).
@@ 4 Methodology
We now outline the general structure of PARCO as shown in Figure 2. First, we formally define parallel multi-agent MDPs for CO (Section 4.1), which we use as a basis to derive our overall parallel autoregressive approach (Section 4.2). We then describe in detail the components of our model: Multi-Agent Encoder (Section 4.3), Communication Layers (Section 4.4), Decoder with Multiple Pointer Mechanism (Section 4.5) and Conflict Handlers (Section 4.6). Finally, we outline the training scheme (Section 4.7).
@@
4.1 Parallel Multi-Agent MDPs
We reformulate the multi-agent MDPs of Section 3.1 as parallel multi-agent MDPs by selecting multiple actions from a joint action space simultaneously to enhance efficiency and coordination (Yu et al., 2022). At each step $t$, ${M}$ agents select actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{M})$ according to a policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}_{1}\times...\times\mathcal{A}_{M}$, which maps the state space $\mathcal{S}$ to the joint action space of agents. A conflict handling function $\psi:(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})$ ensures action compatibility by allowing only one agent to execute when multiple agents select mutually conflicting actions (e.g., the same customer location), and assigning a fallback action (e.g., staying at current position) to the others. Given the resolved agent actions $\tilde{\bm{a}}_{t}=\psi(\bm{a}_{t})$, the state of the problem $s_{t}$ progresses to $s_{t+1}$ according to the transition function $\tau:\mathcal{S}\times(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow\mathcal{S}$. The agents receive a shared reward $R(\bm{a},\bm{x})$, with $\bm{a}=(\tilde{\bm{a}}_{1},...,\tilde{\bm{a}}_{T})$ the sequence of joint agent actions.
In the following, we refer to the entities corresponding to actions in the MDP formulation of a CO problem (e.g., customer locations in VRPs) as nodes, following the convention of Kool et al. (2019a) and Kwon et al. (2020).
@@ 4.2 Parallel AR Models for CO
Motivated by the nature of multi-agent MDPs, PARCO introduces a Parallel AR model for the policy $\pi_{\theta}$. PARCO constructs feasible solutions by simultaneously generating multiple agent actions based on the current state. We formulate the solution generation process in an encoder-decoder framework similarly to Equation 1:
@@
A benefit of our parallel formulation is that the total number of construction steps $T$ can be substantially lower compared to purely AR methods. While the latter require $\sum_{m=1}^{M}T_{m}$ total actions to construct a solution with $T_{m}$ being the number of steps required by agent $m$ to finish its task, PARCO’s Parallel AR needs only $\max_{m}T_{m}$ steps as agents effectively divide the solution space and act concurrently. This leads to a faster solution construction as illustrated in Figures 1 and 5 and significantly reduced training times.

### v3

# PARCO: Parallel AutoRegressive Models for Multi-Agent Combinatorial Optimization

@@ Abstract
Combinatorial optimization problems involving multiple agents are notoriously challenging due to their NP-hard nature and the necessity for effective agent coordination. Despite advancements in learning-based methods, existing approaches often face critical limitations, including suboptimal agent coordination, poor generalization, and high computational latency. To address these issues, we propose PARCO (Parallel AutoRegressive Combinatorial Optimization), a general reinforcement learning framework designed to construct high-quality solutions for multi-agent combinatorial tasks efficiently. To this end, PARCO integrates three key novel components: (1) transformer-based communication layers to enable effective agent collaboration during parallel solution construction, (2) a multiple pointer mechanism for low-latency, parallel agent decision-making, and (3) priority-based conflict handlers to resolve decision conflicts via learned priorities. We evaluate PARCO in multi-agent vehicle routing and scheduling problems, where our approach outperforms state-of-the-art learning methods, demonstrating strong generalization ability and remarkable computational efficiency. We make our source code publicly available to foster future research: https://github.com/ai4co/parco.
@@ 1 Introduction
Figure 1: PARCO generates better solutions with higher efficiency through parallel decision making.
@@
Combinatorial optimization (CO) problems involve determining an optimal sequence of actions in discrete spaces with several crucial domains, including logistics and supply chain management [91]. Many practical CO problems require a solution to be constructed by coordinating multiple distinct entities (i.e., agents), each with unique characteristics. We call such problems multi-agent CO. This class of problems naturally arises in real-world applications such as coordinated vehicle routing for disaster management [60], manufacturing [37] and last-mile delivery optimization [1], where heterogeneous agents must operate under complex constraints.
@@
CO problems are notoriously hard to solve and cannot generally be solved optimally in polynomial time due to their NP-hardness [36, 79]. While traditional methods such as exact and heuristic methods have been developed to solve a variety of problems [51, 25], these approaches often concentrate on single-agent scenarios and require long execution times. Moreover, multi-agent CO poses additional challenges including additional constraints and different optimization objectives, such as minimizing lateness or the makespan among agents [82, 72].
@@
Recently, Neural Combinatorial Optimization (NCO) has emerged as a powerful alternative by learning efficient neural solvers [4]. In particular, Reinforcement Learning (RL) has shown promise due to its ability to learn directly from interactions with CO problems instead of relying on costly labeled datasets in the shape of optimal solutions, and could even outperform traditional approaches by automatically discovering better (neural) heuristics [3, 49, 40]. Among NCO methods, Autoregressive (AR) models—constructing solutions step-by-step—have garnered attention for their ability to generate solutions for a variety of problems with hard constraints [47, 50, 7]. This capability is crucial for addressing complex problems with multiple constraints, such as heterogeneous capacity [2] or precedences in pickup and delivery [77] and machine scheduling problems [101].
@@
However, despite their promising results in single-agent settings, AR-CO methods pose two challenges that hinder their practical adoption in solving complex multi-agent CO problems. First, existing AR methods for multi-agent CO suffer from poor coordination, resulting in unsatisfactory solution quality and poor generalization across varying problem sizes and agent configurations [108, 84, 104, 63]. Moreover, AR sequence generation is associated with high latency due to each single action (or “token”) depending on each previous one, akin to large language models [52, 21]. This issue becomes increasingly pronounced when dealing with large problem instances involving numerous agents.
@@
This paper introduces PARCO (Parallel AutoRegressive Combinatorial Optimization), a novel learning framework to address multi-agent combinatorial problems effectively via parallel solution construction as illustrated in Fig.˜1. We design specialized transformer-based Communication Layers to enhance coordination among agents during decision-making, enabling collaborative behaviors. Our model leverages a Multiple Pointer Mechanism to efficiently generate actions for various agents simultaneously at each solution construction step. Priority-based Conflict Handlers ensure feasibility and resolve potential conflicts among agents’ decisions based on learned priorities.
@@
  • We evaluate PARCO on multi-agent vehicle routing and scheduling, where we outperform state-of-the-art learning methods in solution quality, generalization, and efficiency.
@@ 2 Related Work
Recent advancements in Neural Combinatorial Optimization (NCO) have shown promising end-to-end solutions for combinatorial optimization problems [4, 94]. NCO has led to the development of a variety of methods for diverse problems, including the incorporation of problem-specific biases [35, 64, 44, 41, 83, 20, 31], bi-level solution pipelines [56, 96, 105], learning-guided search [95, 45, 85, 57, 58, 59, 42], improvement methods [26, 69, 70, 28], effective training algorithms [43, 23, 14, 13, 80, 65, 103], downstream applications [11, 106, 71, 66, 93], and the recent development of end-to-end foundation models [61, 107, 15, 6, 34, 53, 22]. Among such methods, RL-based end-to-end autoregressive (AR) models present several advantages, including eliminating the need for labeled solutions, reducing reliance on handcrafted heuristics, and achieving high efficiency in generating high-quality solutions [47, 49, 50, 5, 8].
@@
While the seminal works in AR-CO methods of Vinyals et al. [89], Kool et al. [47], Kwon et al. [49, 50] propose models that can be used in loose multi-agent settings, such methods cannot be employed directly to model heterogeneous agents with different attributes and constraints. Building on Kool et al. [47], Son et al. [84] and Zheng et al. [104] introduce attention-based policies for multi-agent min-max routing. These models adopt a sequential AR construction strategy, solving for one agent at a time and switching agents only after completing a single-agent solution. While this approach outperforms decentralized methods [10, 76] it remains inherently sequential. In contrast, other multi-agent AR methods determine a location for every agent in each solution construction step, but the different agents select their actions sequentially in either random [102] or learned order [18, 55, 63], making these models still suffer from high generation latency as well as missing inter-agent communication and coordination. Zong et al. [108] propose a multi-agent pickup and delivery model with parallel decoding, using distinct decoders for each agent. However, this approach exhibits limited generalizability due to inflexible fixed decoders for specific agents, lacking a powerful communication mechanism, and conflict resolution handled naively by assigning random precedence to agents, restricting the robustness of the model.
PARCO addresses the shortcomings of previous works by leveraging parallel solution construction for any number of agents efficiently with a Multiple Pointer Mechanism, enhancing coordination via Communication Layers, and solving conflicts in a principled manner via Priority-based Conflict handlers. Finally, unlike previous works, PARCO is a general framework tackling multi-agent CO without restricting to a single class of problems.
@@ 3.1 Markov Decision Processes
Multi-agent CO problems can be formulated as Markov Decision Processes (MDPs) and solved autoregressively using RL [74]. In this framework, a solution $\bm{a}$ to a CO problem instance $\bm{x}$ is represented as a sequence of actions. Actions $a_{t}$ are selected sequentially from the action space $\mathcal{A}$ of size $N$ based on the current state $s_{t}\in\mathcal{S}$, which encodes the problem’s configuration at step $t$. In multi-agent problems, at each step one agent $m\in\mathcal{M}=\{1,\dots,M\}$ selects an action $a^{m}_{t}$ according to a policy $\pi_{\theta}$, usually represented by a $\theta$-parametrized neural network, mapping states to actions. Agents are selected either by some predefined precedence rule as in the sequential planning of Son et al. [84], where agent solutions are constructed one after another, or by the policy itself, in which case $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}\times\mathcal{M}$. Given the agent and its corresponding action, the problem then transitions from state $s_{t}$ to state $s_{t+1}$ according to a transition function $\tau:\mathcal{S}\times\mathcal{A}\times\mathcal{M}\rightarrow\mathcal{S}$. This process reaches the terminal state once it has generated a feasible solution $\bm{a}=(a_{1},...,a_{T})$ for the problem instance $\bm{x}$ in $T$ construction steps. The (sparse) reward $R(\bm{a},\bm{x})$ is usually obtained only in the terminal state and takes the form of the negative of the cost function of the respective CO problem.
@@ 3.2 AR Models for CO
Given the sequential nature of MDPs, autoregressive (AR) models pose a natural choice for the policy $\pi_{\theta}$. AR methods construct a viable solution by sequentially generating actions based on the current state and previously selected actions. Without loss of generality, the process can be represented in an encoder-decoder framework as:
@@
where $\bm{a}_{<t}=(a_{1},\ldots,a_{t-1})$ is the sequence of actions taken prior to $t$ and $\bm{h}=f_{\theta}(\bm{x})$ is an encoding of the problem instance $\bm{x}$ obtained via the encoder network $f$. The decoder $g_{\theta}$ then autoregressively generates the sequence of actions, conditioned on $\bm{h}$ and the previously generated actions. The parameters $\theta$ encompass both the encoder and decoder components, which together define the policy as the mechanism for producing the joint distribution $p_{\theta}(\bm{a}|\bm{x})$. Thus, the RL objective becomes finding the optimal set of parameters $\theta^{*}$ that maximizes the reward function $R$ [48, 49, 43].
@@ 4 Methodology
We now outline the general structure of PARCO as shown in Fig.˜2. First, we formally define parallel multi-agent MDPs for CO (§˜4.1), which we use as a basis to derive our overall parallel autoregressive approach (§˜4.2). We then describe in detail the components of our model: Multi-Agent Encoder (§˜4.3), Communication Layers (§˜4.4), Decoder with Multiple Pointer Mechanism (§˜4.5) and Conflict Handlers (§˜4.6). Finally, we outline the training scheme (§˜4.7).
@@
4.1 Cooperative Multi-Agent MDPs
We reformulate the MDPs of §˜3.1 as cooperative multi-agent MDPs [9], often termed fully cooperative Markov games [73], by selecting multiple actions from a joint action space simultaneously to enhance efficiency and coordination [98, 68]. At each step $t$, M{M} agents select actions $\bm{a}_{t}=(a_{t}^{1},...,a_{t}^{M})$ according to a policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}_{1}\times...\times\mathcal{A}_{M}$, which maps the state space $\mathcal{S}$ to the joint action space of agents. A conflict handling function $\psi:(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})$ ensures action compatibility by allowing only one agent to execute when multiple agents select mutually conflicting actions (e.g., the same customer location), and assigning a fallback action (e.g., staying at current position) to the others. Given the resolved agent actions $\tilde{\bm{a}}_{t}=\psi(\bm{a}_{t})$, the state of the problem $s_{t}$ progresses to $s_{t+1}$ according to the transition function $\tau:\mathcal{S}\times(\mathcal{A}_{1}\times...\times\mathcal{A}_{M})\rightarrow\mathcal{S}$. The agents receive a shared reward $R(\bm{a},\bm{x})$, with $\bm{a}=(\tilde{\bm{a}}_{1},...,\tilde{\bm{a}}_{T})$ the sequence of joint agent actions.
In the following, we refer to the entities corresponding to actions in the MDP formulation of a CO problem (e.g., customer locations in VRPs) as nodes, following the convention of Kool et al. [47] and Kwon et al. [49].
@@ 4.2 Parallel AR Models for CO
Motivated by the nature of multi-agent MDPs, PARCO introduces a Parallel AR model for the policy $\pi_{\theta}$. PARCO constructs feasible solutions by simultaneously generating multiple agent actions based on the current state. We formulate the solution generation process in an encoder-decoder framework similarly to Eq.˜1:
@@
A benefit of our parallel formulation is that the total number of construction steps $T$ can be substantially lower compared to purely AR methods. While the latter require $\sum_{m=1}^{M}T_{m}$ total actions to construct a solution with $T_{m}$ being the number of steps required by agent $m$ to finish its task, PARCO’s Parallel AR needs only $\max_{m}T_{m}$ steps as agents effectively divide the solution space and act concurrently. This leads to a faster solution construction as illustrated in Figs.˜1 and 6 and significantly reduced training times.